# SAM-guided Graph Cut for 3D Instance Segmentation

### v2

@@ 5.2 Comparisons with the state-of-the-art methods
Table 4: Analyses of the choice for pseudo-labels generation.
5.3 Ablation studies
We conduct ablation studies on ScanNet to show the effectiveness of each component and design. We evaluate with five configurations: (1) We directly use the edge weight of SAM annotated graph (without GNN) to apply graph cut. (2) We annotate the graph without node features. (3) We annotate the graph without edge weights. (4) We train GNN without regularization loss. (5) Our full method. We report the quantitative results in Tab. 3.
The results indicate that the absence of a GNN leads to a decline in segmentation performance. The most common issue is that some objects or structures are incorrectly connected together, as demonstrated by the examples in Fig. 5. Node features have a certain impact on the predictive performance of the GNN, while edge weights have a very significant impact, almost playing a dominant role. Additionally, we found that regularization loss is also very important.
Furthermore, we conducted experiments to analyse the advantages of using CropFormer to generate pseudo-labels compared to SAM. For the labels generated by CropFormer, we also used SAM to predict a label. The accuracy of both methods was evaluated using the ground truth annotations provided by ScanNet. As indicated in the Tab. 4, CropFormer achieved higher accuracy.
6 Conclusion
In this paper, we introduced a novel 3D segmentation method with SAM guided graph cut. The key idea is to pre-segment 3D scenes into superpoints, and then utilize the prompt mechanism of SAM to assess the affinity scores between superpoints. We propose a GNN based graph cut method to improve the robustness, which is trained with pseudo-labels generated by 2D segmentation network. Experiments showed that the proposed method was able to achieve accurate segmentation results and can generalize well to different datasets.
Figure 5: Ablation studies of GNN. We show some examples demonstrating incorrect segmentations that occurred due to the absence of GNN.
Discussion.
@@
Our method not only requires geometric data (mesh/point cloud) but also needs multi-view images as input, which to some extent limits its application scenarios. Moreover, we perform segmentation based on merging SuperPoint. When an object is part of a SuperPoint, we are unable to segment it out. This situation occurs occasionally, for example, a poster adhered to a wall. To address this, a more sophisticated pre-segmentation model, which consider not only geometric information but also semantics, should be designed. One viable approach is to also use the guidance from SAM or other visual models during the pre-segmentation stage. We leave it to future works.
Appendix
@@
[…]
SAM-based Graph + NCuts 5.1 13.6 39.5 7.5 16.5 34.2 10.1 18.4 31.0
SAM-based Graph + DBSCAN 6.6 14.1 24.7 8.2 15.0 22.6 12.9 23.5 35.6
SAM-based Graph + Graph partition 12.9 30.0 57.7 12.6 24.5 40.2 12.6 25.5 41.9
SAM-based Graph + GNN + NCuts 8.9 21.5 47.7 8.0 16.9 32.4 8.9 16.4 29.4
SAM-based Graph + GNN + DBSCAN 7.1 15.0 26.4 8.3 15.3 23.4 13.3 23.9 36.0
SAM-based Graph + GNN + Graph partition 15.1 33.3 59.1 12.9 25.3 43.6 14.7 28.0 43.2
Table 5: Ablation studies of different graph cut methods.
Appendix A Points sampling in projection mask
To sample $k=5$ points uniformly and not too far from the boundary of the projection mask of each superpoint in each view, we first use the Euclidean Distance Transform to compute the distance from each pixel within the mask to its boundary, creating a distance map. We then select the point with the maximum value in the distance map to ensure it is near the center. To prevent subsequent sampled points from being too close to this first point, we set the values in the distance map within a certain area around this point to zero. This process is iteratively repeated for sampling the remaining points.
Appendix B Structure of GNN
Appendix C Comparison with Panoptic Lifting
Appendix D Analyses of different graph cut method
Appendix E Discussions of SAM guidance
As shown in the ablation studies in our paper, both the node features and edge weights calculated based on SAM are effective for our method, with the edge weights being particularly crucial. To further analyse their effectiveness, we attempted to remove both and use PointNet++ to compute node features. Specifically, we utilized PointNet++ to extract features from the point cloud, averaging the features within a superpoint to serve as the node feature. We employed the same loss function as in our method and optimized the network parameters of both PointNet++ and the GNN simultaneously. We found that this approach resulted in very poor performance.

### v3

@@ 6 Conclusion
Discussion.  Our method not only requires geometric data (mesh/point cloud) but also needs multi-view images as input, which to some extent limits its application scenarios. Moreover, we perform segmentation based on merging superpoints. When an object is part of a superpoint, we are unable to segment it out. This situation occurs occasionally, for example, a poster adhered to a wall. To address this, a more sophisticated pre-segmentation model, which consider not only geometric information but also semantics, should be designed. One viable approach is to also use the guidance from SAM or other visual models during the pre-segmentation stage. We leave it to future works.
Acknowledgement.  The authors would like to acknowledge the support from NSFC (No. 62322207), Ant Group and Information Technology Center and State Key Lab of CAD&CG, Zhejiang University.
Appendix 0.A Superpoint generation
For the mesh of each scene on ScanNet and ScanNet++, we use the segmentator [2] provided by ScanNet, which adopts the algorithm described in [16] based on the mesh’s normals. Specifically, for ScanNet, we directly use the over-segmentation provided in their dataset, which is obtained by running the segmentator with KThresh set to $0.01$ and segMinVerts set to $20$. For ScanNet++, due to the higher point cloud density of ScanNet++, we adjust the parameters to KThresh as $0.2$ and segMinVerts as $500$ to run the segmentator. For KITTI-360, we use the unsupervised point cloud segmentation algorithm proposed in [19], which first computes geometric features for each point from its 3D position and color values and then obtains the partitioned superpoints by minimizing the global energy function, and we set the regularization strength $\rho$ to $0.1$.
Appendix 0.B Points sampling in projection mask
To sample $k=5$ points uniformly and not too far from the boundary of the projection mask of each superpoint in each view, we first use the Euclidean Distance Transform implemented by OpenCV [1] to compute the distance from each pixel within the mask to its boundary, creating a distance map. We then select the point with the maximum value in the distance map to ensure it is near the center. To prevent subsequent sampled points from being too close to this first point, we set the values in the distance map within a certain area around this point to zero. This process is iteratively repeated for sampling the remaining points.
Appendix 0.C Multi-scale mask selection
The 2D segmentation model SAM is designed to output masks at three different scales, each with a corresponding confidence score, to enable segmentation at different granularities. Please refer to their paper for detailed information. In our pipeline, we tend to choose masks with larger areas, as they are more likely to correspond to the segmentation of a complete object. However, when the confidence of a larger area segmentation is high, we will consider masks of smaller areas. Specifically, if the confidence of the largest mask is higher than the others, or lower but within a margin of $0.05$, then we choose this mask. If this criterion is not met, we select from the remaining two masks: if the medium-sized mask has a higher confidence than the smallest mask, or its confidence is lower but within a margin of $0.05$, then we choose the medium-sized mask; otherwise, we choose the smallest mask.
Appendix 0.D Structure of GNN
Appendix 0.E Implementation details of graph cut
When performing segmentation, for every two vertices, if their affinity is below a certain threshold, we consider them to be unconnected. Conversely, if their affinity is above this threshold, to further improve robustness, we identify paths of length 2 between these two vertices. We then record the number of paths where both edges have high affinity scores and the number where one edge is high and the other low. If the ratio of the latter exceeds a predefined threshold, we regard the two vertices to be unconnected; otherwise, they are considered as connected. Once the connection status of each edge is determined using this method, we employ a union-find algorithm [52] to merge all connected superpoints, resulting in the 3D segmentation of the scene.
Appendix 0.F Details of evaluation protocol
We evaluate the class-agnostic AP scores of all methods across all datasets, meaning that during the evaluation, we only consider the accuracy of the masks, without taking into account their semantic categories. This approach follows that of [46]. Since our primary focus is on object instance segmentation, and the baseline Mask3D cannot segment the floors and walls, we exclude predictions of floor and wall regions (only the floor for KITTI-360) from evaluation for all methods to ensure a more fair comparison. Additionally, we also exclude segmentations predicted in other unlabeled regions.
@@ Appendix 0.F Details of evaluation protocol
Appendix 0.G Comparison with Panoptic Lifting
Appendix 0.H Analyses of different graph cut method
Table 7: Ablation studies of different graph cut methods.
@@ Appendix 0.H Analyses of different graph cut method
mAP AP50 AP25 mAP AP50 AP25 mAP AP50 AP25
NCuts 15.7 31.7 59.0 10.1 18.6 34.7 19.5 30.2 45.0
DBSCAN 10.3 18.6 27.8 10.5 17.2 25.0 20.5 31.4 42.1
Graph partition 19.7 37.7 61.6 13.7 25.2 43.0 22.6 36.2 48.5
GNN + NCuts 18.0 35.0 59.4 11.3 20.1 35.2 18.1 27.7 40.7
GNN + DBSCAN 11.0 19.6 29.2 10.7 17.5 25.9 21.1 32.2 43.0
GNN + Graph partition 22.1 41.7 62.8 15.3 27.2 44.3 23.8 37.2 49.1
Appendix 0.I Discussions of SAM guidance
As shown in the ablation studies in our paper, both the node features and edge weights calculated based on SAM are effective for our method, with the edge weights being particularly crucial. To further analyze their effectiveness, we attempted to remove both and use PointNet++ to compute node features. Specifically, we utilized PointNet++ to extract features from the point cloud, averaging the features within a superpoint to serve as the node feature. We employed the same loss function as in our method and optimized the network parameters of both PointNet++ and the GNN simultaneously. We found that this approach resulted in very poor performance.
Appendix 0.J Performance on small objects
We show segmentation results on small objects in Fig. 6. Our method successfully segments a considerable number of small objects, such as several shoes in Fig. 6 and some vases and small items on the shelf in Fig. 6, as framed in red, though some small objects are not separated from each other or their surroundings, as framed in blue.
Figure 6: Segmentation performance on small objects.
Appendix 0.K Failure cases
To better understand the performance of our method, we show two typical failure cases (over-segmention and under-segmention) of our method in Fig. 7, framed in red and blue respectively.
Figure 7: Failure cases.